OpenStudy (linn99123):

The temperature of an oven went from 320° to 280°. What was the percent decrease in temperature? A. 1.25% B. 12.5% C. 1.43% D. 14.3%

OpenStudy (anonymous):

Do you know the percent increase/ decrease formula?

OpenStudy (linn99123):

no

OpenStudy (anonymous):

12.5%

OpenStudy (anonymous):

320-280=40 40/320=0.125 0.125*100=12.5